2007-04-12 ALMA Calibration Group Telecon Agenda and Minutes


ALMA Calibration Group Telecon
  • Date: April 12, 2007 (Thursday)
  • Time: 14:00 UT
  • Duration: 1.0 hr
  • USA Toll Free Number: 1-877-919-7148
  • International: +1-203-566-1039
  • Passcode: 5 1 0 4 6 8
  • Leader: JeffMangum

Table of Contents:



Complete List of Action Items


Action Items from Last Telecon

Assigned To Due Date Description State Notify  


Agenda and Minutes

NOTE: Discussion from the telecon is shown in italic.


Attendees

  1. RobertLucas
  2. JeffMangum
  3. BojanNicholic
  4. AlisonPeck
  5. JohnRicher
  6. AlWootten


Calibration Examples Status

Still waiting on final versions of bandpass (RobertLucas), delay (RobertLucas), polarization (EdFomalont), optics (BaltasarVilaVilaro), and ACA calibration issues (EdFomalont) calibration examples. See CalExamples wiki for further information.


Calibration Sequence Development

Will begin process of turning CalExamples into calibration "sequences" during 2007Q2. Need to first meet with Computing IPT to figure out which Computing IPT personnel will be available to participate in sequence development.


Amplitude Calibration Subsystem Development Reports

  • Berkeley Absolute Flux Calibration System (Jack Welch)
  • Semi-Transparent Vane (STV) Calibration System
    • Need only to get final report from Jesus Martin-Pintado. This is still pending.
  • Multi/Dual-Load (M/DL) Calibration System (Matt Carter)


Weather (Ancilliary) Measurement Devices

In December 2006 JeffMangum submitted weather station instrumentation (RH, P, T, and wind) RFQ documentation to NRAO procurement system. After having reviewed information received from bidders and consulted with Calibration Group (see CalAncillary for discussion), on 2007/03/21 JeffMangum submitted purchase information to RobertLaing for procurement.


Oxygen Sounder

Progress on device characterization reported by JohnRicher. Has developed a contact with one of the companies that produces O2 sounders. Will investigate issues related to operating such a device at 5000 m, among other things.


Allocation of Weather Devices for the OSF

AlWootten notes that the Science IPT needs to provide a plan, with AIV, for weather station installation at the OSF within the next month or two. At the AOS near the TB in the ESO trailer is a Davis weather station. There is another weather station on a tower near the APEX site, according to Nyman. Neither is currently being used. These could be moved to the holography tower at the OSF.

There are two CR4 hygrometers in the US. One has been repaired by the manufacturer (Buck Research) and was shipped back to Chile last week. Another, at the ATF, is nonfunctional. The repaired instrument will be installed at the OSF.

Antenna IPT has requested OSF weather station installed by 5 May. Science IPT has provided all of the elements of this except for one, recently revealed as necessary, solar flux (SF) measurement device. One SF which is on Mangum's CV bookcase. There is one installed at the ATF, a newer one, which Nick bought when he couldn't find the one which is here. Either of these could be moved to Chile but given the export/import rules it would have to happen in a hurry as I understand things are tied up for three weeks when they arrive, and it would take a week to send it. It must be sent through official channels.

The one remaining thing I can check on is how these interface. Glendenning still has our draft ICD; I suspect he wants a computer running linux which could run the python script used at the ATF to get the output from the CR4 through its serial interface. A laptop in the trailer at the AOS ran a perl script to collect the CR4 output. The laptop is in Tucson but reportedly in very bad shape. A computer must be located for the CR4.

So detailed schedule:

  • SF from ATF to SCO? Sramek?
  • SF from SCO to OSF? Nick?
  • CR4 to SCO done 5 Mar 2007
  • CR4 to OSF done?
  • Vaisala windbird to AOS done 2005
  • Vaisala windbird to OSF Who? When?

Once these are at the OSF, these things are in Rick's domain for AIV at the holography tower.

Repair of the 225 GHz tipper was discussed with Donoso. In a short while it could be installed at the AOS and monitored over the internet--this includes leaving it where it is and incorporating it into the AOS infrastructure under operations control.


Solar Observations and Calibration

It has proven difficult to keep the nutator both light enough to nutate and cool enough in the face of the expected solar flux. We have been asked if the four nutators will be necessary for solar observations. Bastian thinks not and is thinking on it. We may need to ensure that these nutators never observe the Sun.

MarkHoldaway has written a draft document on Calibration of Solar Observations. We discussed this document at our last telecon. Few revisions are needed. MarkHoldaway will update and submit final version within the coming few weeks.


Date of Next Phone Meeting

The next ALMA Calibration Group telecon will be 2007, May 10th 14:00 UT.
-- JeffMangum - 11 Apr 2007
Topic revision: r4 - 2007-04-12, JeffMangum
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback